# MagicDriveDiT: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

### v1

@@ 4.3 Progressive Bootstrap Training
This training strategy is grounded in two observations. First, in the controllable generation, we note that the model initially optimizes for higher content quality before learning controllability, a pattern also observed by Gao et al. [12]. Training the model from scratch requires numerous iterations for convergence. Our progressive transition approach enables the model to acquire controllability more rapidly. Second, during the stage transition, we find that the model adapts more quickly to high resolution [5] compared to long-video controllability. Therefore, we train the controllability from the first stage and focus on optimizing more iterations with (short) videos rather than images.
Figure 5: Progressive Bootstrap Training in MagicDriveDiT. For high-resolution long video generation, we train the model to progressively scale up from both resolution and frame count.
4.4 Variable Length and Resolution Adaptation
As stated in Section 4.3, we employ a three-stage training process. In the last two stages, we incorporate videos of varying resolutions and lengths for training. Specifically, in the second stage, we train with a mixture of videos up to 65 frames in length and a maximum resolution of 424$\times$800. In the third stage, we mix videos up to 241 frames (the maximum frame count of the dataset) and a resolution of up to 848$\times$1600 (the maximum resolution of dataset). Details of the mixing method are provided in the Appendix B.
Compared to training with a single resolution and length, this mixed approach allows the model to rapidly comprehend information across the dimensions of resolution and frame count (Section 5.3 holds more ablation study for this). Consequently, our final model supports the generation of various resolutions (224$\times$400, 424$\times$800, 848$\times$1600) and frame counts (1-241 frames). It also enables extrapolation across both dimensions, allowing sampling beyond the training configurations, such as 848$\times$1600 at 129 frames or 424$\times$800 at 241 frames (as shown in Section 6).
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset [4], a prominent dataset for street-view generation [11, 41, 27], to evaluate MagicDriveDiT. We adhere to the official splits, using 700 multi-view videos for training and 150 for validation. Our primary baseline is MagicDrive [11], comprising three models: MagicDrive (16f) for 16-frame video generation; MagicDrive (60f), which extends MagicDrive (16f) to 60 frames, both from Gao et al. [11]; and MagicDrive3D, a 16-frame model from Gao et al. [10]. Further details are provided in Appendix C.
@@
Evaluation Metrics. We evaluate both realism and controllability in street-view generation for videos and images. For video generation, we follow benchmarks from [29], utilizing FVD for video quality; mAP from 3D object detection and mIoU from BEV segmentation for controllability, using BEVFormer [21] for both tasks, which is a video-based perception model. For image generation, we adopt the metrics from Gao et al. [11], employing FID for image quality; mAP with BEVFusion [26] and road mIoU with CVT [49] for controllability assessment. BEVFusion and CVT are image-based models. Our MagicDriveDiT is trained solely on the training set of the nuScenes dataset, and we generate corresponding videos/images for the validation set annotations, applying the aforementioned metrics with perception models pre-trained on real data for evaluation.
Model Setup. In accordance with Section 4.2, we adopt the 3D VAE framework from CogVideoX [43] and train diffusion models for street-view video generation from scratch [^3] For more discussion on this choice, see Section 5.3 and Appendix F.. Initially, the model comprises only spatial blocks for base and control blocks, focusing on image generation. Temporal blocks are incorporated in the second stage, forming the full architecture of MagicDriveDiT. Sequence parallelism is employed during the final stage to manage GPU memory limitations, detailed in Appendix A. Further training details can be found in Appendix D.
@@ 5.2 Results and Analysis
Generation Quality. MagicDriveDiT excels in both video and image generation tasks. In video tasks, it significantly reduces FVD compared to MagicDrive (Table 2), due to the DiT architecture enhancing inter-frame consistency and spatial-temporal condition encoding for precise control of object motion and positioning. Further illustrated in Figure 6, MagicDriveDiT generates high-resolution videos that not only improve quality but also incorporate more intricate details, closely resembling footage captured by real cameras. This enhancement is achieved through our advanced training on variable lengths and resolutions, allowing for more realistic and detailed outputs.
Also beneficial from the mixing training approach, MagicDriveDiT is capable of image generation. As shown in Table 3, MagicDriveDiT matches baseline performance in multi-view street-view tasks and surpasses baselines in vehicle segmentation mIoU and object detection mAP. This demonstrates the strong generalization capabilities of our spatial-temporal condition encoding.
Controllability. The quantitative results presented in Table 2 and 3 demonstrate that the images and videos generated by MagicDriveDiT effectively reflect the control conditions. Additionally, Figure 8 provides visual results, showing that multiple control conditions can independently influence the generated content. For instance, weather can be altered via text input (from sunny to rainy) while maintaining road structure, and the trajectories of other vehicles and the ego vehicle. By varying the combinations of conditions, MagicDriveDiT is capable of producing a diverse range of high-quality street-view videos.
@@
Figure 7: Reconstruction Visualization from Different VAEs. CogVAE [43] maintains most details compared with others and exhibits better performance for high-resolution contents.
Figure 8: MagicDriveDiT generates high-resolution (e.g., 424$\times$800) street-view videos for 241 frames (i.e., the full length of nuScenes videos, approximately 20 seconds at 12 FPS) with multiple controls (i.e., road map, object boxes, ego trajectory, and text). Notably, the 241-frame length at 424$\times$800 are unseen during training, demonstrating our method’s generalization capability to video length. We annotate the ego-vehicle trajectory and selected objects to aid localization, with same-color boxes denoting the same object. Due to space constraints, the “rainy” example includes only two frames; additional examples can be found in the Appendix H.
@@ 5.3 Ablation Studies
VAE Comparison for Street Views. Before training the diffusion models, we evaluate the performance of open-source 3D VAEs (i.e., CogVAE [43] and Open-Sora [47]) on street views, in comparison with the 2D SD VAE [31]. As illustrated in Figure 7, CogVAE consistently outperforms its counterparts on reconstruction ability. Besides, as presented in Appendix E, CogVAE demonstrates minimal performance degradation over longer video sequences, making it particularly well-suited for long video generation tasks. Additionally, we found that all VAEs exhibit improved reconstruction capabilities with increasing resolution. This insight is beneficial for advancing our model’s ability to generate high-quality images and videos by focusing on higher-resolution outputs. More detailed quantitative results and discussion see Appendix E.
@@
Figure 9: Validation Loss through Training with Different SP Encodings. 4×4$\times$4 ×down (our methods in MagicDriveDiT) can help the model converge, performing the best among all the encodings.
@@
Progressive Bootstrap Training. The three-stage progressive training approach markedly improves model training efficiency relative to direct Stage 3 training. Table 4 indicates that, over 4 days, for example, Stage 1 executes approximately 60 times more iterations than Stage 3, and Stage 2 achieves about 7 times more iterations. The progressive training is vital for controlled generative models, which require extensive iterations for effective convergence, as discussed in Section 4.3. The progressive strategy enables the rapid acquisition of high-quality video generation capabilities, utilizing faster iterations in the early stages to enhance convergence and expedite learning.
Stages Sec./Iter. Iter. for 4 days
stage 1 4.32 80k
stage 2 39.84 8.7k
stage 3 264.96 1.3k
Table 4: Speed for Each Training Stage of MagicDriveDiT, measured on NVIDA A800 GPUs. Over a 4-day period (for example), Stage 1 training yields nearly 60 times more iterations than Stage 3, and Stage 2 offers about 7 times more. This value is calculated by multiplication with sequence parallel (SP) size (in practice, we use SP size of 4 for the stage 3, with 66.24s/it).
Variable Length and Resolution Training. As presented in Section 4.4, MagicDriveDiT adopts a training strategy by mixing different length and resolution videos. Our ablation study demonstrates the effectiveness of this strategy. As shown in Table 5, the limitations of VAE (as discussed in Section 5.2) are evident when using only low-resolution videos (17$\times$224$\times$400), indicated by worse FVD and lower mAP and mIoU scores compared to other settings. Incorporating longer videos improves the model’s controllability (both mAP and mIoU are higher), while incorporating high-resolution videos enhances overall quality (all three metrics are notably improved)
@@ 6 Applications
Extrapolation for Longer Video Generation. Through variable length and resolution training (Section 4.4), MagicDriveDiT effectively generates videos exceeding the training setup’s length. Though trained on videos up to 129$\times$424$\times$800 and 241$\times$224$\times$400, MagicDriveDiT successfully produces 241$\times$424$\times$800 videos (full length of nuScenes), which are double the length of the training samples, as demonstrated in Figure 8. To further validate this capability, we compare the first-16-frame FVD for short videos (17 frames) and per-16-frame FVD for longer videos (65 frames and above). As shown in Table 6, the 16-frame FVDs remain consistent across both seen and extrapolated configurations, extending to 129$\times$848$\times$1600$\times$6 generation (this is not the upper limit; please see note in the Appendix). This affirms the model’s robust generalization capability.
@@
Table 6: Generation Quality for Videos Longer than Training. We randomly sample 10 sequence from the nuScenes validation set and report FVD (the lower the better). $n\times$: $n$ times of maximum training frame number, i.e., 129 frames for 424$\times$800 and 33 for 848$\times$1600. /: exceed the maximum frame of dataset.
@@ 7 Conclusion and Discussion
In this paper, we presented MagicDriveDiT, an innovative framework for high-resolution and long video synthesis with precise control, specifically tailored for applications like autonomous driving. By leveraging the DiT architecture and Flow Matching, MagicDriveDiT effectively addresses the challenges of scalability and geometry control in video generation. Our approach introduces novel spatial-temporal conditional encoding and a progressive bootstrapping strategy with variable length and resolution adaption, enhancing the model’s generalization ability. Extensive experiments demonstrate that MagicDriveDiT generates realistic videos that maintain spatial and temporal coherence, significantly surpassing previous methods in resolution and frame count. This work advances video synthesis and opens new possibilities for simulations and other applications in autonomous driving.
@@
Note: Our model is capable of generating videos at a resolution of 848$\times$1600 for 241 frames, which is the highest resolution and frame count in the nuScenes dataset. However, the inference cost is currently substantial. Therefore, the primary numerical results in our paper do not utilize this maximum setting, nor do we emphasize this capability in our main paper. We have included some generated results in Appendix H for reference. Future work may focus on further reducing the inference cost.
@@ B More Details for Mixed Resolution and Frames Training
MagicDriveDiT is trained through progressive bootstrap approach (see Section 4.3) with variable length and resolution data configurations (see Section 4.4). Consequently, our method of data mixing corresponds to the three training stages, as detailed in Table I.
@@
C More Experimental Details
The nuScenes dataset includes 12Hz unannotated data and 2Hz annotated data. According to our experiments, high-frame-rate videos are more beneficial for generative model learning. Therefore, we follow [11] and interpolated the 2Hz annotations to 12Hz annotations with ASAP [36]. Although the interpolation results are not entirely accurate, they do not affect the training for video generation.
Semantic Classes for Generation. We follow  [11] in data setup. Specifically, for objects, ten categories include car, bus, truck, trailer, motorcycle, bicycle, construction vehicle, pedestrian, barrier, and traffic cone. For the road map, eight categories include drivable area, pedestrian crossing, walkway, stop line, car parking area, road divider, lane divider, and roadblock.
D More Training Details
Optimization. We train our diffusion models using Adam optimizer and a constant learning rate at $8e^{-5}$ with a 3000-step linear warm-up in the last two stages. We primarily use 32 NVIDA A800 GPUs (80G) for training. Our model can also trained with Ascend 910B (64G). The batch size for each stage is set according to the iteration speed, following the bucket strategy as [47]. For example, in stage 2, we set the batch size for 33$\times$424$\times$800 to 1, which takes about 30s/it. Then we set the batch size to other video types to achieve about 30s/it. This strategy can ensure the load balance between different GPU processes.
@@ D More Training Details
E More Comparison among VAEs
To quantitatively compare the performance of the VAE, we randomly selected two 6-view videos from the nuScenes dataset and used the PSNR metric to evaluate the VAE’s reconstruction ability. Table II presents the results, averaged across the six views. From these results, we observe that CogVAE [43] demonstrates the best reconstruction ability, even surpassing the 2D VAE [31]. Comparing the results from different settings, we find that the current 3D VAEs exhibit good generalization ability for long videos, primarily due to the window-based downsampling techniques [43, 47]. Additionally, we observe that high-resolution content retains a relatively high PSNR after VAE reconstruction, indicating that the current VAEs are more favorable for high-resolution data. This observation also supports our motivation for high-resolution generation.

### v2

@@ 4.5 Progressive Training with Mixed Video Types
To accelerate model convergence, we employ a three-stage training strategy: starting with low-resolution images, transitioning to high-resolution short videos, and finally using high-resolution long videos. This approach is based on two key insights. First, in controllable generation, models prioritize content quality before learning controllability, also observed by Gao et al. [12]. Our progressive training enables faster controllability acquisition. Second, models adapt more quickly to high resolution than to long videos, so we emphasize short-video training early on. Details of the mixing method are provided in the Appendix B.
In the third stage, we mix videos up to 241 frames (the maximum frame count of the dataset) at 224$\times$400 and a resolution of up to 848$\times$1600 (the maximum resolution of nuScenes [4]) at 33 frames. This mixed-resolution and mixed-length training allows the model to generalize across dimensions, supporting flexible generation and extrapolation beyond training configurations (e.g., up to 848×1600 at 241 frames as demonstrated in Section 6). Ablation studies in Section 5.3 further validate this approach.
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset [4], a prominent dataset for street-view generation [11, 40, 26], to evaluate MagicDriveDiT. We adhere to the official splits, using 700 multi-view videos for training and 150 for validation. Our primary baseline is MagicDrive [11], comprising three models: MagicDrive (16f) for 16-frame video generation; MagicDrive (60f), which extends MagicDrive (16f) to 60 frames, both from Gao et al. [11]; and MagicDrive3D, a 16-frame model from Gao et al. [10]. Moreover, we employed the Waymo Open Dataset [33] to further fine-tune the model, aiming to validate its scalability and to facilitate the generation of videos with diverse environmental styles and different numbers of perspectives. Further details are provided in Appendix D.
@@
Evaluation Metrics. We evaluate both realism and controllability in street-view generation for videos and images. For video generation, we follow benchmarks from [28], utilizing FVD for video quality; mAP from 3D object detection and mIoU from BEV segmentation for controllability, using BEVFormer [20] for both tasks, which is a video-based perception model. For image generation, we adopt the metrics from Gao et al. [11], employing FID for image quality; mAP with BEVFusion [25] and road mIoU with CVT [48] for controllability assessment. BEVFusion and CVT are image-based models.
Model Setup. In accordance with Section 4.2, we adopt the 3D VAE framework from CogVideoX [42] and train diffusion models for street-view video generation from scratch [^3] For more discussion on this choice, see Section 5.3 and Appendix H.. Initially, the model comprises only spatial blocks for base and control blocks, focusing on image generation. Temporal blocks are incorporated in the second stage, forming the full architecture of MagicDriveDiT. Further training details can be found in Appendix E.
@@ 5.2 Results and Analysis
Generation Quality. MagicDriveDiT excels in both video and image generation tasks. In video tasks, it significantly reduces FVD compared to MagicDrive (Table 2), due to the DiT architecture enhancing inter-frame consistency and spatial-temporal condition encoding for precise control of object motion and positioning. Further illustrated in Figure 6, MagicDriveDiT generates high-resolution videos that not only improve quality but also incorporate more intricate details, closely resembling footage captured by real cameras. This enhancement is achieved through our advanced training on variable lengths and resolutions, allowing for more realistic and detailed outputs. We further conduct human evaluation on consistency in Appendix F.
Beneficial from the mixing training approach, MagicDriveDiT is capable of image generation. As shown in Table 3, MagicDriveDiT matches baseline performance in multi-view street-view tasks and surpasses baselines in vehicle segmentation mIoU and object detection mAP. This demonstrates the strong generalization capabilities of our spatial-temporal condition encoding.
Controllability. The quantitative results presented in Table 2 and 3 demonstrate that the images and videos generated by MagicDriveDiT effectively reflect the control conditions. Additionally, Figure 5 and 8 provide visual results, showing that multiple control conditions can independently influence the generated content. For instance, weather can be altered via text input (from sunny to rainy) while maintaining road structure, and the trajectories of other vehicles and the ego vehicle. By varying the combinations of conditions, MagicDriveDiT is capable of producing a diverse range of high-quality street-view videos.
@@
Figure 7: Reconstruction Visualization from Different VAEs. CogVAE [42] maintains most details compared with others and exhibits better performance for high-resolution content.
Figure 8: MagicDriveDiT generates high-resolution (e.g., 424$\times$800 here) street-view videos for 241 frames (i.e., the full length of nuScenes videos, approximately 20 seconds at 12 FPS) with multiple controls (i.e., road map, object boxes, ego trajectory, and text). Notably, the 241-frame length at 424$\times$800 is unseen during training, demonstrating our method’s generalization capability to video length. We annotate the ego-vehicle trajectory and selected objects to aid localization, with same-color boxes denoting the same object. Due to space constraints, the “rainy” example includes only two frames; additional examples can be found in Appendix J and on our project page.
@@ 5.3 Ablation Studies
VAE Comparison for Street Views. Before training the diffusion models, we evaluate the performance of open-source 3D VAEs (i.e., CogVAE [42] and Open-Sora [46]) on street views, in comparison with the 2D SD VAE [30]. As illustrated in Figure 7, CogVAE consistently outperforms its counterparts on reconstruction ability. Besides, as presented in Appendix G, CogVAE demonstrates minimal performance degradation over longer video sequences, making it particularly well-suited for long video generation tasks. Additionally, we found that all VAEs exhibit improved reconstruction capabilities with increasing resolution. This insight is beneficial for advancing our model’s ability to generate high-quality images and videos by focusing on higher-resolution outputs. For more detailed quantitative results and discussion, see Appendix G.
@@
Figure 9: Validation Loss through Training with Different ST Encodings. 4×4$\times$4 ×down (our methods in MagicDriveDiT) can help the model converge, performing the best among all the encodings.
@@
Variable Length and Resolution Training. As presented in Section 4.5, MagicDriveDiT adopts a training strategy by mixing different length and resolution videos. Our ablation study demonstrates the effectiveness of this strategy. As shown in Table 4, the limitations of VAE (as discussed in Section 5.2) are evident when using only low-resolution videos (17$\times$224$\times$400), indicated by worse FVD and lower mAP and mIoU scores compared to other settings. Incorporating longer videos improves the model’s controllability (both mAP and mIoU are higher), while incorporating high-resolution videos enhances overall quality (all three metrics are notably improved).
@@ 6 Applications
Extrapolation for Longer Video Generation. Through variable length and resolution training (Section 4.5), MagicDriveDiT effectively generates videos exceeding the training setup’s length. Though trained on videos up to 33$\times$848$\times$1600 and 241$\times$224$\times$400, MagicDriveDiT successfully produces 241$\times$848$\times$1600 videos (full length and full resolution of nuScenes), which are 8×8$\times$8 × the length of the training samples, as demonstrated in Figure 8 and videos on our project page. To further validate this capability, we compare the first-16-frame FVD for short videos (17 frames) and per-16-frame FVD for longer videos (65 frames and above). As shown in Table 5, the 16-frame FVDs remain consistent across both seen and extrapolated configurations, extending to 129$\times$848$\times$1600$\times$6 generation (this is not the upper limit; please see note in the Appendix). This affirms the model’s robust generalization capability.
@@
Table 5: Generation Quality for Videos Longer than Training. We randomly sample 10 sequences from the nuScenes validation set and report FVD (the lower the better). $n\times$: $n$ times of maximum training frame number, i.e., 129 frames for 424$\times$800 and 33 for 848$\times$1600. /: exceed the maximum frame of dataset.
Fast Generalization on Other Datasets. To evaluate the generalization of MagicDriveDiT, we fine-tuned the Stage 3 model on the Waymo dataset [33], enabling rapid generation of 3-view videos in $1$ day ($1$k+ steps) with strong controllability (examples shown in Figure 5 and more videos can be found on our project page). Moreover, mixed training on Waymo and nuScenes datasets further enhanced the model’s ability to generate videos with varying perspectives, improving overall quality. The final model achieved an FVD of 105.17 on Waymo and 74.30 on nuScenes, surpassing results in Table 2.
@@ 7 Conclusion and Discussion
In this paper, we presented MagicDriveDiT, an innovative framework for high-resolution and long video synthesis with precise control, specifically tailored for applications like autonomous driving. By introducing MVDiT block and spatial-temporal conditional encoding module, MagicDriveDiT effectively addresses the challenges of scalability and geometry control in video generation with the DiT architecture and 3D VAE. Our approach also introduces richer contextual description for driving video generation, and a progressive training strategy with variable length and resolution adaption, enhancing the model’s generation ability. Extensive experiments demonstrate that MagicDriveDiT generates realistic videos that maintain spatial and temporal coherence, significantly surpassing previous methods in resolution and frame count, opening new possibilities for simulations and other applications in autonomous driving.
@@
Note: Our model is capable of generating videos at a resolution of 848$\times$1600 for 241 frames, which is the highest resolution and frame count in the nuScenes dataset. However, the inference cost is currently substantial. Therefore, the primary numerical results in our paper do not utilize this maximum setting. We have included some generated results in Appendix J and on our project website for reference. Future work may focus on further reducing the inference cost.
@@ B More Details for Mixed Resolution and Frames Training
MagicDriveDiT is trained through a progressive training approach with variable length and resolution data configurations (see Section 4.5). Consequently, our method of data mixing corresponds to the three training stages, as detailed in Table I.
@@
C Efficiency of Progressive Bootstrap Training
The three-stage progressive training approach markedly improves model training efficiency relative to direct Stage 3 training. Table II indicates that, over 4 days, for example, Stage 1 executes approximately 60 times more iterations than Stage 3, and Stage 2 achieves about 7 times more iterations. The progressive training is vital for controlled generative models, which require extensive iterations for effective convergence, as discussed in Section 4.5. The progressive strategy enables the rapid acquisition of high-quality video generation capabilities, utilizing faster iterations in the early stages to enhance convergence and expedite learning.
Stages Seconds/Iter. Iter. for 4 days
stage 1 4.32 80k
stage 2 39.84 8.7k
stage 3 264.96 1.3k
Table II: Speed for Each Training Stage of MagicDriveDiT, measured on NVIDA A800 GPUs. Over a 4-day period (for example), Stage 1 training yields nearly 60 times more iterations than Stage 3, and Stage 2 offers about 7 times more. This value is calculated by multiplication with sequence parallel (SP) size (in practice, we use SP size of 4 for the stage 3, with 66.24s/it).
D More Experimental Details
The nuScenes [4] dataset includes 12Hz unannotated data and 2Hz annotated data. According to our experiments, high-frame-rate videos are more beneficial for generative model learning. Therefore, we follow [11] and interpolated the 2Hz annotations to 12Hz annotations with ASAP [36]. Although the interpolation results are not entirely accurate, they do not affect the training for video generation. The Waymo Open Dataset [33] includes 10Hz annotated data. We follow the official splitting, which has 798 clips for training and 202 clips for validation. The original dataset contains 5 views. However, the dimensions of the left and right side views are smaller than those of the three front views, and their field of view is limited. Therefore, we only retain three front views for training and validation.
Semantic Classes for Generation. We follow  [11] in data setup on nuScenes. Specifically, for objects, ten categories include car, bus, truck, trailer, motorcycle, bicycle, construction vehicle, pedestrian, barrier, and traffic cone. For the road map, eight categories include drivable area, pedestrian crossing, walkway, stop line, car parking area, road divider, lane divider, and roadblock. For Waymo, the semantics for objects include pedestrian, car, and cyclist, while the semantics for road maps include drivable area, crosswalk, road line (yellow), and road line (white).
E More Training Details
Optimization. We train our diffusion models using Adam optimizer and a constant learning rate at $8e^{-5}$ with a 3000-step linear warm-up in the last two stages. We primarily use 32 NVIDA A800 GPUs (80G) for training. Our model can also be trained with Ascend 910B (64G). The batch size for each stage is set according to the iteration speed, following the bucket strategy as [46]. For example, in stage 2, we set the batch size for 33$\times$424$\times$800 to 1, which takes about 30s/it. Then we set the batch size to other video types to achieve about 30s/it. This strategy can ensure the load balance between different GPU processes.
@@ E More Training Details
F Human Evaluation for Multi-frame & Multi-view Consistency
Evaluating multi-frame and multi-view consistency in video generation has long been a challenging issue [1], as the academic community lacks a unified standard to accurately assess such consistency. To address this, we employed human evaluation to measure these two aspects of consistency. Specifically, we invited participants with diverse backgrounds to compare videos generated by MagicDriveDiT and Gao et al. [11] under identical conditions. Participants were asked to select the video with better consistency, and the winning probability of each model was statistically analyzed. As illustrated in Figure II, MagicDriveDiT demonstrated significantly superior consistency, indicating a substantial improvement over Gao et al. [11].
Figure II: Experts evaluated content consistency against Gao et al. [11], showing MagicDriveDiT generates videos with superior consistency.
G More Comparison among VAEs
To quantitatively compare the performance of the VAE, we randomly selected two 6-view videos from the nuScenes dataset and used the PSNR metric to evaluate the VAE’s reconstruction ability. Table III presents the results, averaged across the six views. From these results, we observe that CogVAE [42] demonstrates the best reconstruction ability, even surpassing the 2D VAE [30]. Comparing the results from different settings, we find that the current 3D VAEs exhibit good generalization ability for long videos, primarily due to the window-based downsampling techniques [42, 46]. Additionally, we observe that high-resolution content retains a relatively high PSNR after VAE reconstruction, indicating that the current VAEs are more favorable for high-resolution data. This observation also supports our motivation for high-resolution generation.